OpenStudy (anonymous):

Solve the inequality and type the solution in interval notation. No spaces! -25 ≤ -3x - 4<-13

OpenStudy (anonymous):

-25 ≤ -3x - 4<-13 -21 ≤ -3x < -9 9 ≤ x < 3

OpenStudy (anonymous):

$$-25\le-3x-4\lt-13\\-21\le-3x\lt-9\text{ by adding 4 to each "side"}\\7\ge x\gt3\text{ or, equivalently, }3<x\le7$$This may also be expressed in interval notation as \(x\in(3,7]\).